OpenStudy (anonymous):

how to simplify (x-3)-(x+1)(2x+5)

OpenStudy (anonymous):

this is just an expression ?

OpenStudy (anonymous):

expand (x+1)(2x+5) using product rule of polynomials

OpenStudy (anonymous):

x-3-2x^2-2x-5x-5 =-(2x^2+6x+8) =-2(x^2+3x+4)

OpenStudy (anonymous):

By F.O.I.L (First Outer Inner Last) we can expand: \[(x+1)(2x+5) = 2x^2+5x+2x+5=2x^2+7x+5\] Then we've got: \[(x-3)-(2x^2+7x+5)=(2x^2+6x+2)\] ...man, I am super rusty. I still for some reason think I've done something wrong!

OpenStudy (anonymous):

yeah ur wrong..this is the correct answer... -2x^2+8x+2

OpenStudy (anonymous):

Well now we've got three answers...lets try to come to agreement (not that I claim I am certainly correct...attempting again!

OpenStudy (anonymous):

dat's 5n.